# Waveform Design of Multi-User-Multi-Target ISAC System based on Kullback–Leibler Divergence

### v1

@@ II-A Communication System
At each $l$ instance, a data symbol $s_{\mathrm{c},k,l}$ intended for the $k$-th UE is drawn from a certain normalised constellation, i.e, $\mathbb{E}[\left|s_{\mathrm{c},k,l}\right|^{2}]=1$. The received signal at the $k$-th UE can be formulated as follows,
@@
where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, the variable $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ represents the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent, and $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l}$ id the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN). The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$. In the next two subsections, we revisit the conventional MRT and ZF designs which are typically used to precode the communication data symbols at the BS. These precoding schemes will serve as benchmarks for our designs.
@@ II-A1 MRT beamforming
In MRT precoding, the amount of received power at the UEs is maximized without taking into account the imposed interference. This approach usually lead to acceptable performance at low SNR values but its performance deteriorates at large SNRs. The normalized MRT beamforming matrix is designed for all UEs based on the channel matrix as follows,
@@ II-A2 ZF beamforming
On the contrary, ZF precoding eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power. This approach usually results in excellent performance at high SNR values. The normalized ZF beamforming matrix is designed also based on the channel matrix, and can be represented as follows,
@@ II-B Radar system
The radar system proposed in this study is designed for high adaptability, allowing for real-time adjustments in detection capabilities on a per-frame basis, where each frame is composed of $L$ snapshots. A key parameter of this system is $T$, representing the count of radar beams emitted during a specific detection frame. By harnessing MIMO radar technology, multiple beams can be generated simultaneously using orthogonal signals [30, 31]. Consequently, $T$ signifies the upper limit of detectable targets within a single frame.
This radar system’s design is robust, as it can explore multiple angular-range-Doppler bins over successive frames, thereby increasing the chance of identifying more targets. Furthermore, its flexibility in varying the number of emitted beams allows for the simultaneous detection of a larger set of targets, providing the ability to fine-tune the system’s detection capability according to different operational demands. However, it is crucial to acknowledge that the value of $T$ is naturally constrained by the available number of antennas and the total number of UEs being serviced by the BS, highlighting the critical need for efficient waveform design, and resource allocation strategies in ISAC systems.
Our focus aligns with scenarios where targets are spatially separated, ensuring each target is confined to a distinct radar bin, as explored in prior studies [32, 33]. Additionally, the literature offers a range of algorithms developed to separate signals from closely spaced or unresolved targets, facilitating accurate target enumeration [34]. A straightforward method to count detected targets involves applying the detection mechanism across all radar bins spanning the angular, range, and Doppler dimensions. The total radar return signal at the $l$-th snapshot is shown as follows,
@@
where $\mathbf{H}_{t}\in\mathbb{C}^{N\times N}$ is the target response matrix of the BS-Target-BS route for the $t$-th target, following a Rayleigh distribution [35]. The $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the additive white Gaussian noise vector at the $l$-th snapshot. To analyze the radar returns from all beams, a series of matched filters are typically applied, each tuned to a specific waveform $s_{\mathrm{r},t}:\forall t=\left\{1,2,\cdots,T\right\}$, which corresponds to a particular radar angular-range-Doppler bin. Because $s_{\mathrm{r},t}$ is orthogonal to $s_{\mathrm{r},i}$ for all $t\neq i$, the radar returns from distinct beams can be effectively separated, enabling independent detection of each potential target. The radar precoding matrix $\mathbf{W}_{r,l}$ is designed such that the covariance matrix is $\mathbf{R}_{w}\triangleq\frac{1}{L}\sum_{l=1}^{L}\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}\in\mathbb{C}^{N\times N}$. The received signal for the $t$-th target at the BS, under the binary hypotheses $\mathcal{H}_{q}\forall q\in\{0,1\}$, where $\mathcal{H}_{1}$ signifies target presence and $\mathcal{H}_{0}$ denotes target absence, can be mathematically represented as follows,
@@
III The KLD of communication system: ZF and MRT precoding
In this section, we derive the KLD of the communication system for two typical scenarios, namely, the MRT, and the ZF beamforming schemes using the shared deployment antenna configuration. The analysis is carried out for the normalized precoders introduced in Section II-A.
In order to, derive the KLD, it is necessary to establish the statistical properties of the signal, IUI, radar interference, and noise components. For a pair of multivariate Gaussian distributed random variables having mean vectors of $\mathbf{\mu}_{m}$ and $\mathbf{\mu}_{n}$ and covariance matrices of $\Sigma_{m}$ and $\Sigma_{n}$, the KLD can be derived as,
$\mathrm{KLD}_{n\rightarrow m}=\frac{1}{2\ln 2}\left(\mathrm{tr}\left(\Sigma_{n}^{-1}\Sigma_{m}\right)-2+\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)^{T}\right.\\ \;\left.\times\;\Sigma_{n}^{-1}\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)+\ln\frac{\left|\Sigma_{n}\right|}{\left|\Sigma_{m}\right|}\right),$ (7)
The application of this formula to the communication system requires the determination of the mean vectors and covariance matrices for each precoding scheme.
@@ III-A MRT beamforming
For the MRT precoding scheme, the analysis encompasses the desired signal, IUI, radar interference, and noise components. The examination starts with the IUI term, which is the second term in (2), which can be written as,
@@
Since that the elements of $\mathbf{h}_{k}$ and $\mathbf{h}_{i}$ $\forall i\neq k$ are i.i.d. complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$, the term $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ is the inner product of two independent complex Gaussian vectors. Therefore, central limit theorem (CLT) can be invoked to show that the distribution of $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ tends to be complex Gaussian with zero mean and variance $N\sigma_{h}^{4}$ for a significant number of antennas $N$. On the other hand, the normalization term $\|\mathbf{H}\|_{F}^{2}=\sum_{i=1}^{K}\|\mathbf{h}_{i}\|^{2}$ follows a chi-squared distribution with $2NK$ degrees of freedom, scaled by $\sigma_{h}^{2}/2$. Consequently, the exact distribution of $\omega_{\mathrm{MN},k}$ is a ratio of complex normal variables to the square root of a chi-squared variable. Unfortunately, this ratio distribution does not have a simple closed-form expression. However, it is worth noting that we can approximate its behavior for large $N$ and $K$, by using the law of large numbers, $\|\mathbf{H}\|_{F}^{2}\approx NK\sigma_{h}^{2}$. Consequently, using this approximation, we can write (8) as,
@@
Now, $\omega_{\mathrm{MN}}$ is a sum of independent complex normal variables. By invoking the CLT, for large $KN$, this sum approaches a complex normal distribution with zero mean and a variance of,
@@
The accuracy of this approximated density of $\omega_{\mathrm{MN}}$ is validated by simulation and shown in Fig.1, which compares it with simulated density functions under different values of $N$. It can be also shown that the distribution of $\eta_{k}$ can be well approximated by a complex Gaussian distribution using the CLT, particularly for large $N$. This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of,
@@
where $\mathbf{R}_{w,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{\tilde{w}}_{\mathrm{r},t,l}\mathbf{\tilde{w}}_{\mathrm{r},t,l}^{H}$. Now, the equivalent IUI and radar interference plus noise term $\tilde{\omega}_{\mathrm{MN},k}=\omega_{\mathrm{MN}}+\eta_{k}$ can be approximated as a complex Gaussian random variable with zero mean and variance $\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2}$. In order to find the KLD for the MRT case with matrix normalization, we need to evaluate the expected value of the desired signal power, which is represented by the first term in (2). This term, $|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$ can be expressed as follows,
@@
After deriving the statistical properties of the signal and interference components, the KLD formula can be applied. Given that $\Sigma_{n}=\Sigma_{m}=(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})\mathbf{I}_{2}$, and $\mathbf{\mu}_{k,m}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(m)}$, and $\mathbf{\mu}_{k,n}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(n)}$, these can be substituted into equation (7), then the KLD can be found as [21, Corollary 1],
@@
where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ which depends on the transmitted data constellation, and thus $\lambda$ is constant for a given constellation, for example, for MPSK signalling, $\lambda_{\mathrm{MPSK}}=2\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left(1-\cos\left(\phi_{k,m}-\phi_{k,n}\right)\right)$. Finally, substituting (13) in (14), then the KLD for MRT precoding with matrix normalization can be found as,
@@
Figure 1: The density functions of the approximated random variable $\omega_{MN}$.
@@ III-B ZF beamforming
With ZF precoding, the received signal at the $k$-th user can be rewritten as follows,
@@
where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalized ZF precoding matrix. It should be noticed that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder completely eliminates the IUI. The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are shown as follows,
@@
where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalization of the ZF precoding matrix. Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, it can be shown that $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter $N-K+1$ and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$. Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$. By substituting (17), and (18) in (7), then the KLD can be reduced to,
@@
where the expectation of $1/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$ can be calculated using the PDF of Gamma distribution as follows,
@@
Substituting this result into the expression for $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}$ in (19), the final expression for the average KLD in the case of ZF precoder with matrix normalization can be reduced to,

### v2

@@ II-A Communication System
At each $l$ instance, a data symbol $s_{\mathrm{c},k,l}$ intended for the $k$-th UE is drawn from a normalised constellation, i.e, $\mathbb{E}[\left|s_{\mathrm{c},k,l}\right|^{2}]=1$. The received signal at the $k$-th UE can be formulated as
@@
where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ representing the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l}$ is the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN). The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$. In the next two subsections, we revisit the conventional MRT and ZF designs which are typically used to precode the communication data symbols at the BS. These precoding schemes will serve as benchmarks for our designs.
@@ II-A1 MRT beamforming
In MRT precoding, the amount of received power at the UEs is maximised without taking into account the imposed interference. This approach usually leads to acceptable performance at low SNR values but its performance deteriorates at large SNRs. The normalised MRT beamforming matrix is designed for all UEs based on the channel matrix as follows,
@@ II-A2 ZF beamforming
On the contrary, ZF precoding eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power. This approach usually results in excellent performance at high SNR values. The normalised ZF beamforming matrix is designed also based on the channel matrix, and can be represented as follows,
@@ II-B Radar system
The radar system is designed for high adaptability, allowing for real-time adjustments in detection capabilities across $L$ frame snapshots. $T$ represents the radar beams emitted during a specific detection frame. Using MIMO radar technology, multiple beams can be generated simultaneously using orthogonal signals [30, 31], where $T$ signifies the upper limit of detectable targets in a single frame.
The system explores multiple angular-range-Doppler bins over successive frames, increasing target identification. Its flexibility in varying emitted beams allows for the simultaneous detection of more targets, to fine-tune the system’s detection capability according to different operational demands. However, it is crucial to acknowledge that the value of $T$ is constrained by the available antennas and UEs serviced by the BS, highlighting the need for efficient waveform design and resource allocation in ISAC systems.
Our focus aligns with scenarios where targets are spatially separated, with each target confined to a distinct radar bin, as per prior studies [32, 33]. The literature presents algorithms to separate signals from closely spaced or unresolved targets, facilitating accurate target enumeration [34]. The total radar return signal at the $l$-th snapshot is shown as follows,
@@
where $\mathbf{H}_{t}\in\mathbb{C}^{N\times N}$ is the target response matrix of the BS-Target-BS route for the $t$-th target, following a Rayleigh distribution [35]. The $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the additive white Gaussian noise vector at the $l$-th snapshot. To analyse the radar returns from all beams, a series of matched filters are typically applied, each tuned to a specific waveform $s_{\mathrm{r},t}:\forall t=\left\{1,2,\cdots,T\right\}$, which corresponds to a particular radar angular-range-Doppler bin. Because $s_{\mathrm{r},t}$ is orthogonal to $s_{\mathrm{r},i}$ for all $t\neq i$, the radar returns from distinct beams can be effectively separated, enabling independent detection of each potential target. The radar precoding matrix $\mathbf{W}_{r,l}$ is designed such that the covariance matrix is $\mathbf{R}_{w}\triangleq\frac{1}{L}\sum_{l=1}^{L}\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}\in\mathbb{C}^{N\times N}$. The received signal for the $t$-th target at the BS, under the binary hypotheses $\mathcal{H}_{q}\forall q\in\{0,1\}$, where $\mathcal{H}_{1}$ signifies target presence and $\mathcal{H}_{0}$ denotes target absence, can be mathematically represented as follows,
@@
III The communication KLD: ZF and MRT
In this section, we derive the KLD of the communication system for the MRT and ZF beamforming using the shared deployment antenna configuration. The analysis is carried out for the normalised precoders introduced in Section II-A.
To derive the KLD, it is necessary to establish the statistical properties of the signal, IUI, radar interference, and noise. For a pair of multivariate Gaussian distributed random variables having mean vectors of $\mathbf{\mu}_{m}$ and $\mathbf{\mu}_{n}$ and covariance matrices of $\Sigma_{m}$ and $\Sigma_{n}$, the KLD can be derived as,
$\mathrm{KLD}_{n\rightarrow m}=\frac{1}{2\ln 2}\left(\mathrm{tr}\left(\Sigma_{n}^{-1}\Sigma_{m}\right)-2+\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)^{T}\right.\\ \;\left.\times\;\Sigma_{n}^{-1}\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)+\ln\frac{\left|\Sigma_{n}\right|}{\left|\Sigma_{m}\right|}\right).$ (7)
@@ III-A MRT beamforming
For the MRT precoding scheme, the analysis encompasses the desired signal, IUI, radar interference, and noise components. The examination starts with the IUI term, which is the second term in (2), given by,
@@
Since the elements of $\mathbf{h}_{k}$ and $\mathbf{h}_{i}$ $\forall i\neq k$ are i.i.d. complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$, the term $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ is the inner product of two independent complex Gaussian vectors. Thus, the central limit theorem (CLT) can be invoked to show that the distribution of $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ tends to be complex Gaussian with zero mean and variance $N\sigma_{h}^{4}$ for a large number of antennas $N$. Meanwhile, the normalisation term $\|\mathbf{H}\|_{F}^{2}=\sum_{i=1}^{K}\|\mathbf{h}_{i}\|^{2}$ follows a chi-squared distribution with $2NK$ degrees of freedom, scaled by $\sigma_{h}^{2}/2$. Consequently, the exact distribution of $\omega_{\mathrm{MN},k}$ is a ratio of complex normal variables to the square root of a chi-squared variable. This ratio distribution does not have a simple closed-form expression. However, we can approximate its behaviour for large $N$ and $K$, by using the law of large numbers, $\|\mathbf{H}\|_{F}^{2}\approx NK\sigma_{h}^{2}$. Using this approximation, we can write (8) as,
@@
$\omega_{\mathrm{MN}}$ is a sum of independent complex normal variables. By invoking CLT, for large $KN$, the sum approaches a complex normal distribution with zero mean and a variance of
@@
The accuracy of the approximated density of $\omega_{\mathrm{MN}}$ is validated by simulation, which compares it with simulated density functions under different values of $N$. It can also be shown that the distribution of $\eta_{k}$ is well approximated by a complex Gaussian distribution using CLT, particularly for large $N$. This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of
@@
where $\mathbf{R}_{w,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{\tilde{w}}_{\mathrm{r},t,l}\mathbf{\tilde{w}}_{\mathrm{r},t,l}^{H}$. Now, the equivalent IUI and radar interference plus noise term $\tilde{\omega}_{\mathrm{MN},k}=\omega_{\mathrm{MN}}+\eta_{k}$ can be approximated as a complex Gaussian random variable with zero mean and variance $\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2}$. In order to find the KLD for the MRT case with matrix normalisation, we need to evaluate the expected value of the desired signal power, which is represented by the first term in (2). This term, $|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$ can be expressed as follows,
@@
Given that $\Sigma_{n}=\Sigma_{m}=(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})\mathbf{I}_{2}$, and $\mathbf{\mu}_{k,m}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(m)}$, and $\mathbf{\mu}_{k,n}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(n)}$, substituting these into equation (7), then the KLD can be found as [21, Corollary 1],
@@
where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ is constant for a given constellation, for example, for MPSK signalling, $\lambda_{\mathrm{MPSK}}=2\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left(1-\cos\left(\phi_{k,m}-\phi_{k,n}\right)\right)$. Substituting (13) in (14), the KLD for MRT precoding with matrix normalisation is,
@@ III-B ZF beamforming
With ZF precoding, the received signal at the $k$-th user is,
@@
where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalised ZF precoding matrix. Note, that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI. The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are given by,
@@
where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalisation of the ZF precoding matrix. Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter $N-K+1$ and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$. Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$. By substituting (17), and (18) in (7), then the KLD becomes,
@@
The expectation of $1/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$ can be calculated using the PDF of Gamma distribution as follows,
@@
Substituting this result into (19), we get the final $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}$,